# Lily Erickson EmerGen LLC Minneapolis, MN

### v1

@@ Abstract
To do so, we’ve retrofitted our Cross-Axis Transformer with special embedding and logit layers that allow it to process raw byte data - we call it Mage.
@@ 1 Introduction
Mage uses Cross-Axis Attention from our most recently published paper, because it’s more computationally efficient[erickson2023cat] than its peers.
@@ 2.3 The solution
Architecture Type Cross-Axis Transformer
@@ 3 Experiments
If our Mage is able to learn a mathematical representation of an elliptic curve, it’s not going to get there with only 35 million samples. Modern transformers are trained on tens of trillions of tokens, and so our puny hardware is certainly not competing with the big dogs out in the field.
@@ 4 Scaling and Unit Discrepency
And for completeness’ sake, we will include numbers for both our basic baseline Mage, as well as for an open-source, massive language model: Llama 3.1 405B.
@@
Mage inference baseline FLOPs $2.5*10^{10}$
Mage Total Cycles per Keypair $6.84*10^{9}$
@@ 5 Cracking the 256-bit secp256r1 curve
Total cycles to memorize all keypairs (Mage) $1.57*10^{77}*4.75*10^{6}=1.07*10^{87}$
@@ 5.1 The Birthday Paradox
Base Mage Curve Memorization 5 $1.07*10^{87}$
@@ 6 Conclusion
If our small 784M parameter Mage model was capable of memorizing enough secp256r1 keypairs to break 50% of all private keys, it would manage to do so using only the work required to break 100 keypairs the traditional way. And remember, this invalidates the entire curve.
@@
I’m calling the shot right here and now: 256-bit cryptography is dead. It’s only a matter of time before someone puts a super-computer on this problem and then leaks the model weights.
If you work in cyber-security and are not currently worried about AI, you should be. And if you don’t, then you’d better be using a VPN these days.
@@ 7 Ethics
This, coupled with the fact that open-source software is typically only important for reproducibility insofar as the required results are - well, difficult to reproduce; we find it prudent to state that we will not be open-sourcing Mage’s model architecture, learned weights, or training code for this project at this time. Please see the open-source repository for the original Cross-Axis Transformer[erickson2023cat] if you would like a peak at the original novel model architecture, or our Github repository for the experimental results of this project in CSV format.

### v2

@@ 1 Introduction
We’ll use our Cross-Axis Transformer from our most recently published paper, because it’s more computationally efficient[erickson2023cat] than its peers.
@@ 2.3 The solution
Model Cross-Axis Transformer
@@ 3 Experiments
If our CAT is able to learn a mathematical representation of an elliptic curve, it’s not going to get there with only 35 million samples. Modern transformers are trained on tens of trillions of tokens, and so our puny hardware is certainly not competing with the big dogs out in the field.
@@ 4 Scaling and Unit Discrepency
And for completeness’ sake, we will include numbers for both our basic baseline CAT, as well as for an open-source, massive language model: Llama 3.1 405B.
@@
CAT inference baseline FLOPs $2.5*10^{10}$
CAT Total Cycles per Keypair $6.84*10^{9}$
@@ 5 Cracking the 256-bit secp256r1 curve
Total cycles to memorize all keypairs (CAT) $1.57*10^{77}*4.75*10^{6}=1.07*10^{87}$
@@ 5.1 The Birthday Paradox
Base CAT Curve Memorization5 $1.07*10^{87}$
@@ 6 Conclusion
If our small 784M parameter CAT model was capable of memorizing enough secp256r1 keypairs to break 50% of all private keys, it would manage to do so using only the work required to break 100 keypairs the traditional way. And remember, this invalidates the entire curve.
@@
To add fuel to the fire, even when compressing the public key, storing a raw mapping of 1 trillion keypairs ($(33+32)*8*10^{12}=2.89^{14}$) would have taken just under 300 terabytes of data. Previously, this meant that our particular attack vector here was a read/write (I/O) problem. It would have been impossible to sample such a table in real time. Here however, you simply need a single forward pass through the AI and you’ve instantly got your solution. This is a novel brute-force channel over the entire algorithm.
Realistically, classical encryption needs orders of magnitude more complexity to survive. Simply doubling the key-length is no longer a sufficient deterrent.
Long-term, we need solutions in which there is not a perfect mapping between public and private key.
And with the groundwork laid, I’m willing to call the shot right here and now: classical 256-bit cryptography is dead. It’s only a matter of time before someone puts a super-computer on this problem and then leaks the model weights.
If you work in cyber-security and are not currently worried about AI, you should be. And if don’t, then you’d better be using a VPN these days.
@@ 7 Ethics
This, coupled with the fact that open-source software is typically only important for reproducibility insofar as the required results are - well, difficult to reproduce; we find it prudent to state that we will not be open-sourcing our model architecture, learned weights, or training code for this project at this time. Please see the open-source repository for the original Cross-Axis Transformer[erickson2023cat] if you would like a peak at the original novel model architecture, or our Github repository for the experimental results of this project in CSV format.